ROVIBRONIC ENERGY TOPOGRAPHY I: TENSOR EIGENVALUE STRUCTURE AND TUNNELING EF-FECTS IN LOW-SYMMETRY SPECIES-CLUSTERS OF HIGH SYMMETRY MOLECULES

\underline{JUSTIN C. MITCHELL}, WILLIAM G. HARTER, Department of Physics, University of Arkansas, Fayetteville, AR 72701.

Centrifugal and Coriolis effects on rovibrational eigensolutions may be analyzed and approximated by rotational-energy-surfaces (RES). This technique is extended to study high rank tensor Hamiltonians that model spectra of symmetric molecules such as CH_{4}, SF_{6} and C_{60}. Of particular interest here is superfine structure, which was first studied^{a} with a fourth order rotational Hamiltonian and experimentally confirmed for SF_{6}^{b} and CH_{4}^{c}. The theoretical framework for this study is reviewed for higher angular momentum and higher order Hamiltonians. Earlier studies show symmetry reduction into clusters of local C_{4} and C_{3} symmetry. Here we study effects of combinations of fourth, sixth and higher rank tensor Hamiltonians to investigate a reduction to C_{2} and C_{1} (no symmetry) clusters that involve tunneling between 12 or 24 equivalent local phase-space regions. An understanding of this low symmetry tunneling may prove useful in analysis of internal rotations and provide examples of monodromy.

aJ.P. Aldridge et. al J Mol Spec 58, 165 (1975)

bW.G. Harter, C.W. Patterson, Phys Rev Lett, 38, 244 (1977)

cA. Pines, J. Opt Soc 66, 97 (1976)